Some students planned for a get-together. The budget for food was $500. Five of the students failed to come because of the distance and therefore the cost of food for each member increased by $5. How many students attended the get-together?
This is what I have...Let x=original cost per student. Then 500/x is the original number of students. Five students failed to come. So first get the expression for the number of students that came from this information. Also, cost for each student that came increased by $5. So the x+5=new cost per student; and 500/(x+5) is the number of students that came. Set the two expressions that represent the number of that came equal and solve for x.

Answer by stanbon(75887)

EQUATION:
original average = new average + $5
500/x = 500/(x+5) +5
100/x = 100/(x+5) + 1
100(x+5)= 100x +x(x+5)
100x+500 = 100x + x^2+5x
x^2+5x-500=0
(x+25)(x-20)=0
x=20 (original number of students)
Cheers,
Stan H.